# Efficient RL Training for Reasoning Models via Length-Aware Optimization

### v1

@@ 5.1 Component Ablation
As shown in Figure 5, our proposed reward designs significantly improve both response length control and validation accuracy. The left plot shows that the Standard baseline generates overly long responses, while the Kimi baseline collapses to very short outputs. In contrast, our designs (D1, D1+D2, D1+D3) progressively stabilize length generation, with Short-RL achieving the most balanced outcome. The right plot illustrates consistent accuracy gains from our designs. Short-RL consistently achieving the highest accuracy. Even partial configurations (D1+D2, D1+D3) outperform Kimi, underscoring the effectiveness and complementarity of each design component.
@@
5.2 Impact of Length and Accuracy Tolerance
@@
6 Track the Length Reward
@@
Figure 7: Tracking the length reward during training
@@
We track the proposed metrics and the average response length during training in two experiments, as shown in Figure 7. The curves for SimpleRL-Reason and Open-Reasoner-Zero can be found in Appendix 8.1.3.
7 Conclusions
In this article, we propose a simple yet effective method that can be seamlessly integrated into the on-policy reinforcement learning process for long-form reasoning models. Our approach reduces response length during training while maintaining — and in some cases even improving — overall performance.

### v2

@@ Introduction
As shown in Figure 3, our proposed reward designs significantly improve both response length control and validation accuracy. The subfigure 3a shows that the Standard baseline generates overly long responses, while the Kimi baseline collapses to very short outputs. In contrast, our designs (D1, D1+D2, D1+D3) progressively stabilize length generation, with Short-RL achieving the most balanced outcome. The subfigure 3a illustrates consistent accuracy gains from our designs. Short-RL consistently achieving the highest accuracy. Even partial configurations (D1+D2, D1+D3) outperform Kimi, underscoring the effectiveness and complementarity of each design component. Note that the ppl5 test set here differs from the final evaluation set, following the practice of Logic-RL.
@@
Impact of Length and Accuracy Tolerance
@@
Track the Length Reward
@@
Figure 5: Visualization of the length control rate during training. The plot shows how the length reward is distributed throughout the training process.
@@
, We track the proposed metrics and the average response length during training in two experiments, as shown in Figure 5. We observe that the length reward is distributed throughout the training process. In DeepScaleR, length rewards are applied more frequently. The curves for SimpleRL-Reason and Open-Reasoner-Zero can be found in Appendix Track the Length Reward.
Conclusions
In this article, we propose a simple yet effective method that can be seamlessly integrated into the on-policy reinforcement learning process for long-form reasoning models. Our approach reduces response length during training while maintaining — and in some cases even improving — overall performance. Future work can explore adaptive optimization of the hyperparameters introduced in reward designs II and III.
Limitations
Our method is designed for tasks where responses consist of a reasoning process followed by a short definitive answer (e.g., math, logic). In such settings, lengthy reasoning often contains redundant steps, making efficiency improvements viable. However, for tasks like creative writing, where reasoning is minimal or stylistic variation is valuable, favoring shorter outputs may not be appropriate.
Moreover, reward designs II and III rely on manual hyperparameter tuning, which may require adaptation across different tasks or models.
Appendix
Setting Logic-RL DeepScaleR Open Reasoner Zero SimpleRL-Reason
learning rate 1e-6 1e-6 5e-7 5e-7
batch size 8 128 64 16
ppo_mini_batch_size 32 64 256 64
ppo_micro_batch_size 8 32 64 2
rollout_n 8 8 8 8
temperature 0.7 0.6 1.0 1.0
kl_loss_coef 0.001 0.001 0.001 0.0001
epochs 3 3 1 3
max_response_length 4096 8192 4096 8192
algorithm reinforce++ grpo grpo grpo
$\tau_{\ell}$ 200 100 100 50
$\tau_{\text{acc}}$ 0.05 0.05 0.02 0.05
$\alpha$ 1 1 1 1
Model Qwen2.5-7B DeepSeek Distill Qwen-1.5B Qwen2.5-7B Qwen2.5-7B
Table 4: Training details.
Additional Experiments
(a) Open Reasoner Zero
(b) SimpleRL-Reason
Figure 6: Visualization of the length control rate during train- ing.
Track the Length Reward
We also track the metric defined in Section Track the Length Reward in Figure 6 (Open Reasoner Zero and SimpleRL-Reason).
Figure 7: The prompt template for Logic-RL and Math-RL.
Training Details
Our experiments were conducted using a compute node equipped with 8 NVIDIA H100 GPUs. The CUDA version we use is 12.3.
Logic-RL Training and Evaluation Details
The training and evaluation prompt template (Figure 7a) used in Logic-RL remains the same as in the original GitHub project. The training hyperparameters are listed in Table 4. During evaluation, we directly use the code from Logic-RL, which applies a temperature of 1.0 and top_p=1.0 for logic tasks, and a temperature of 0.8 with top_p= 0.95 for math tasks.
Training and Evaluation Details for Math
The training and evaluation prompt template for three math settings is shown in Figure 7b. The training hyperparameters are listed in Table 4. During evaluation, we directly use the code from DeepScaleR, which employs a temperature of 1.0.
Reward Details
In all the math experiments, the standard reward employs a format and outcome-based reward scheme. That is:
[…]
. In Logic-RL experiments, we directly use their original standard reward design.